A Complete Bibliography of Publications in Marine Ecology Progress Series: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

Abortion [LR79]. Abundance [Sim79]. Accumulation [Klo79]. Acids [BEJ^{+}79, Sie79]. Action [VD79]. Acute [TSF79]. Aggression [She79]. Along [BFC79]. Amino [BEJ^{+}79, Sie79]. Ammonia [RM79]. Ammonium [RM79]. Amphiascoides [Wal79]. Anaerobic [Sch79c]. andromeda [Neu79c]. Annelida [Neu79a, Neu79b, Sch79c]. Apparent [VD79]. Ascidian [FM79]. Ascidians [RR79]. Asterias [Sie79]. Autoradiographic [Rie79]. Axiothella [Wei79].

Bacteria [Rie79, Sie79]. Bacterial [Neu79c]. Balanophyllia [Ger79]. Bank [Sim79]. Barrier [Sim79]. bemhardus [Daw79]. Benthic [Obr79, Spa79]. Benthos [Thi79]. Between [ON79, She79]. Biology [HH79]. Blennius [DV79, VD79]. Brachyuran [SPv79]. Breeding [RL79b]. Byssogenesis [RM79].

Cadmium [Klo79, TSF79]. Calanus [Sch79b]. California [BFC79]. californicus [BFC79]. Capanulinidae [HH79]. Carbohydrates [BEJ^{+}79]. Caribbean [Sim79]. Cassiopea [Neu79c]. Causes [PW79]. Central [BFC79]. Changes [BEJ^{+}79]. Character [Fri79a, Fri79b]. Chemical [MRBLS79]. Cnidaria [Neu79c]. Coast [BFC79]. Coastal [ON79]. Coefficients [Spa79]. Colonial [HH79]. Colonizing [Obr79]. Competition [Fri79b]. Considerations [TB79]. Conversion [Neu79b]. Copepod [Wal79]. Copepoda [BFC79]. Coral [Ger79, RL79a, RL79b]. Corals [LR79, She79]. Crab [SPv79]. Critical [Ger79]. Crustacea [Spa79, Fri79a, Fri79b].

Data [Thi79]. Decapoda [Daw79]. Deep [Thi79]. Dense [Wei79]. Density [Wal79]. Density-Dependent [Wal79]. Dependent [Wal79]. Deposit [Wei79]. Deposit-Feeding [Wei79]. Depth [Lun79]. Determinants [Wei79]. Determined [Lan79]. Development [BEJ^{+}79, Daw79, Fon79]. diadema [Klo79]. Different [Lun79, Neu79b]. Displacement [Fri79a, Fri79b]. Dissolved [BEJ^{+}79, Sie79]. Distribution [DAHL79, She79, Sim79]. Distributions [Wei79]. Diuretic [KKS79]. Drag [Spa79]. Dynamic [VD79].

Ecological [Wei79]. Effect [KKS79, LR79]. Effects [Daw79, FM79, Neu79a, RM79, TSF79]. Efficiency [RR79]. Eggs [Fon79]. Eirene [HH79]. elegans [Ger79]. Elutriation [TB79]. Embryos [RI79]. Equatorial [Ger79]. Evidence [Fri79a]. Excluding [Ger79]. Excreted [Lan79]. Excretion [Lan79].

Factors [Neu79a]. Feeding [Sch79b, Wei79]. Field [Fri79a]. Filtration [FM79, RR79]. First [Thi79]. Fish [DV79, VD79]. Fishes [ON79]. Fishing [ON79]. Flatfishes [PW79]. Fluctuating [DV79]. Food [Neu79b, VD79]. Foraminifer [SHB^{+}79]. Four [RR79]. Fractions [Sim79]. Free [ZM79]. Free-Living [ZM79].

Gastropod [Sch79a]. Genetics [BFC79]. Gland [KKS79]. Glands [RI79]. Gonads [RL79a]. Gravity [Spa79]. Gross [Lan79]. Ground [ON79]. Growth [Fon79, HH79, Lun79, Neu79b].

Harpacticoida [BFC79, Wal79]. Hastigerina [SHB^{+}79]. Hatching [RI79]. Helgoland [Lun79]. helgolandicus [Sch79b]. Herring [RI79]. Heterotrophic [Lan79, Sie79]. Hydrodynamic [Spa79]. Hydroids [HH79]. Hydrozoa [TSF79].

II [BEJ^{+}79, Fri79b, RL79b]. III [HH79]. Improvements [TB79]. Induced [LR79]. Induction [Neu79c]. Influence [Fon79, Neu79b, HH79]. Inhabiting [Lun79]. Interrelationships [ON79]. Interspecific [She79]. Invertebrates [Obr79]. Isopoda [Fri79a, Fri79b].

Japan [ON79].

Kinetics [Lan79].

Laboratory [Fon79]. Laminaria [Lun79]. Laomedea [TSF79]. Larvae [Fon79, Neu79c, SPv79]. Larval [Daw79, Obr79]. Leptomedusa. [HH79]. Life [Ger79, WVDL79]. Living [ZM79]. Locomotory [SPv79]. Loop [KKS79]. loveni [TSF79]. Lunar [SHB^{+}79].

Macro [Rie79]. Macro-Autoradiographic [Rie79]. mammillat [FM79]. Marine [Lan79, Obr79, Spa79, Wal79, WVDL79, ZM79]. Matter [Sim79]. Mechanisms [Wal79]. Metabolism [Sch79c]. Metamorphosis [Neu79c]. Methodological [TB79]. Microbiological [MRBLS79]. Microscale [DAHL79]. Migration [SPv79]. Mixtures [Sch79b].

Natural [BEJ^{+}79, Lan79]. near [Sim79]. Nematodes [ZM79]. Nereis [Neu79a, Neu79b, Sch79c]. North [Lan79, Lun79, Rie79]. Nutrient [BEJ^{+}79]. Nutrient-Poor [BEJ^{+}79]. Nutrition [HH79]. Nutritional [Sie79].

Observations [Fon79]. Offshore [WVDL79]. Oil [LR79, WVDL79]. Ophryotrocha [Klo79]. Outdoor [BEJ^{+}79]. Oxygen [FM79].

Paguridae [Daw79]. Pagurus [Daw79]. Papillomas [PW79]. Parameters [MRBLS79]. Particle [RR79]. Particulate [Sim79]. Patches [Wei79]. Pelagic [ON79]. pelagica [SHB^{+}79]. Periodicity [SHB^{+}79]. Perna [RM79]. Phaeophyceae [Lun79]. Phallusia [FM79]. pholis [DV79, VD79]. Phytoplankton [DAHL79, Lan79, Sch79b]. pistillata [RL79a, RL79b]. Plankton [BEJ^{+}79, ON79]. Planktonic [SHB^{+}79]. Planula [Neu79c]. Planulae [RL79a, RL79b]. Platform [WVDL79]. Pollution [LR79]. Polychaeta [Klo79, Neu79a, Neu79b]. Polychaete [Wei79]. Poor [BEJ^{+}79]. Population [BEJ^{+}79, BFC79, Neu79a, Wal79, BFC79]. Production [ZM79]. Products [Lan79]. Properties [Spa79]. Prosobranchia [Sch79a]. Pumping [FM79].

Quantitative [Thi79].

Rate [RR79]. Rates [Lan79]. Reassessment [Sie79]. Rectal [KKS79]. Red [RL79a, RL79b, Thi79]. Reef [She79, Sim79]. Reference [She79, Sie79]. Region [Lun79]. Regulation [Wal79]. Relation [DAHL79]. Reproduction [RL79a, RL79b, SHB^{+}79]. Responses [DV79]. Retention [RR79]. Rhizostomeae [Neu79c]. Rhythm [SPv79]. Role [Sie79]. rubens [Sie79]. rubrocincta [Wei79].

Salinities [DV79]. Salinity [Daw79, Fon79, Neu79b, TSF79]. Salt [KKS79]. Scyphozoa [Neu79c]. Sea [BEJ^{+}79, Lun79, Sie79, Lan79, ON79, Rie79, RL79a, RL79b, Thi79]. Seasonality [RL79b]. Seasonally [Neu79a]. Secretion [KKS79]. Sediment [Rie79]. Serpulorbis [Sch79a]. Sessile [Sch79a]. Settlement [Neu79c]. Shallow [Ger79]. Shark [KKS79]. Shedding [RL79b]. Short [Lan79, MRBLS79]. Short-Term [Lan79, MRBLS79]. Significance [Sie79, SPv79]. Size [Sim79]. Skin [PW79]. Solea [Fon79]. Solitary [Ger79]. South [Ger79]. Southern [Lan79]. Southwestern [ON79]. sp [Wal79]. Space [Fri79b]. Special [Sie79]. Species [Ger79, Lun79, Neu79b, RR79, Sch79c]. Specific [VD79, ZM79, Spa79]. Sperm [Sch79a]. Sphaeroma [Fri79a, Fri79b]. Spionid [Wei79]. spp [Fri79a, Fri79b]. Stage [Ger79]. Star [Sie79]. Strategies [Lun79, Obr79]. Structure [BFC79]. Studies [Rie79]. Stylophora [RL79a, RL79b]. Sublittoral [Lun79]. Submergence [Ger79]. succinea [Neu79a]. Synchronization [RL79b].

Tank [BEJ^{+}79]. Temperature [Daw79, Fon79, HH79, Neu79b, TSF79]. Tension [FM79]. Term [Lan79, MRBLS79]. Thecata [HH79]. Thecata-Leptomedusa. [HH79]. Their [PW79, She79]. Theoretical [TB79]. Thermocline [DAHL79]. Three [Lun79, PW79, Sch79c]. Tigriopus [BFC79]. Tolerance [RM79]. Toxicity [TSF79]. Transfer [Sch79a]. Transintegumentary [Sie79]. Types [PW79].

Uptake [FM79, Klo79, Lan79, Sie79].

Variations [MRBLS79]. Varying [Neu79a]. Vermetidae [Sch79a]. Vertical [DAHL79, SPv79]. viridis [RM79]. viridula [HH79].

Water [BEJ^{+}79, Ger79]. Within [Wei79].

Zones [Lun79].

References

Brockmann:1979:DNP

[BEJ^{+}79] U. H. Brockmann, K. Eberlein, H. D. Junge, E. Maier-Reimer, and D. Siebers. The development of a natural plankton population in an outdoor tank with nutrient-poor sea water. II. Changes in dissolved carbohydrates and amino acids. Marine Ecology Progress Series, 1:283–291, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

           articles/meps/1/m001p283.pdf.

Burton:1979:PGT

[BFC79] R. S. Burton, M. W. Feldman, and J. W. Curtsinger. Population genetics of Tigriopus californicus (Copepoda: Harpacticoida): I. Population structure along the Central California Coast. Marine Ecology Progress Series, 1:29–39, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p029.pdf.

Derenbach:1979:VMD

[DAHL79] J. B. Derenbach, H. Astheimer, H. P. Hansen, and H. Leach. Vertical microscale distribution of phytoplankton in relation to the thermocline. Marine Ecology Progress Series, 1:187–193, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p187.pdf.

Dawirs:1979:ETS

[Daw79] R. R. Dawirs. Effects of temperature and salinity on larval development of Pagurus bemhardus (Decapoda, Paguridae). Marine Ecology Progress Series, 1:323–329, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p323.pdf.

Davenport:1979:RFB

[DV79] J. Davenport and O. Vahl. Responses of the fish Blennius pholis to fluctuating salinities. Marine Ecology Progress Series, 1:101– 107, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p101.pdf.

Fiala-Medioni:1979:EOT

[FM79] A. Fiala-Médioni. Effects of oxygen tension on pumping, filtration and oxygen uptake in the ascidian Phallusia mammillat. Marine Ecology Progress Series, 1:49–53, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p049.pdf.

Fonds:1979:LOI

[Fon79] M. Fonds. Laboratory observations on the influence of temperature and salinity on development of the eggs and growth of the larvae of Solea solea. Marine Ecology Progress Series, 1:91–99, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p091.pdf.

Frier:1979:CDSa

[Fri79a] J. O. Frier. Character displacement in Sphaeroma spp. (Isopoda: Crustacea). I. Field evidence. Marine Ecology Progress Series, 1: 159–163, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p159.pdf.

Frier:1979:CDSb

[Fri79b] J. O. Frier. Character displacement in Sphaeroma spp. (Isopoda: Crustacea). II. Competition for space. Marine Ecology Progress Series, 1:165–168, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/1/m001p165.pdf.

Gerrodette:1979:ESS

[Ger79] T. Gerrodette. Equatorial submergence in a solitary coral, Balanophyllia elegans, and the critical life stage excluding the species from shallow water in the south. Marine Ecology Progress Series, 1: 227–235, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p227.pdf.

Hundgen:1979:BCH

[HH79] M. Hundgen and M. Hartmann. The biology of colonial hydroids. III. Influence of temperature and nutrition on growth of Eirene viridula (Thecata-Leptomedusa. Capanulinidae). Marine Ecology Progress Series, 1:223–226, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/1/m001p223.pdf.

Kinne:1979:ELD

[KKS79] R. Kinne and E. Kinne-Saffran. Effect of ‘loop diuretics’ on the salt secretion in shark rectal gland. Marine Ecology Progress Series, 1: 129–132, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p129.pdf.

Klockner:1979:UAC

[Klo79] K. Klockner. Uptake and accumulation of cadmium by Ophryotrocha diadema (Polychaeta). Marine Ecology Progress Series, 1: 71–76, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p071.pdf.

Lancelot:1979:GER

[Lan79] C. Lancelot. Gross excretion rates of natural marine phytoplankton and heterotrophic uptake of excreted products in the Southern North Sea, as determined by short-term kinetics. Marine Ecology Progress Series, 1:179–186, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/1/m001p179.pdf.

Loya:1979:AEC

[LR79] Y. Loya and B. Rinkevich. Abortion effect in corals induced by oil pollution. Marine Ecology Progress Series, 1:77–80, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p077.pdf.

Luning:1979:GST

[Lun79] K. Luning. Growth strategies of three Laminaria species (Phaeophyceae) inhabiting different depth zones in the sublittoral region of Helgoland (North Sea). Marine Ecology Progress Series, 1:195– 207, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p195.pdf.

Meyer-Reil:1979:STV

[MRBLS79] L.-A. Meyer-Reil, M. Bolter, G. Liebezeit, and W. Schramm. Shortterm variations in microbiological and chemical parameters. Marine Ecology Progress Series, 1:1–6, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p001.pdf.

Neuhoff:1979:ESV

[Neu79a] H.-G. Neuhoff. Effects of seasonally varying factors on a Nereis succinea population (Polychaeta, Annelida). Marine Ecology Progress Series, 1:263–268, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/1/m001p263.pdf.

Neuhoff:1979:ITS

[Neu79b] H.-G. Neuhoff. Influence of temperature and salinity on food conversion and growth of different Nereis species (Polychaeta, Annelida). Marine Ecology Progress Series, 1:255–262, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p255.pdf.

Neumann:1979:BIS

[Neu79c] R. Neumann. Bacterial induction of settlement and metamorphosis in the planula larvae of Cassiopea andromeda (Cnidaria: Scyphozoa, Rhizostomeae). Marine Ecology Progress Series, 1:21– 28, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p021.pdf.

Obrebski:1979:LCS

[Obr79] S. Obrebski. Larval colonizing strategies in marine benthic invertebrates. Marine Ecology Progress Series, 1:293–300, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p293.pdf.

Ogawa:1979:IBP

[ON79] Y. Ogawa and T. Nakahara. Interrelationships between pelagic fishes and plankton in the coastal fishing ground of the southwestern Japan Sea. Marine Ecology Progress Series, 1:115–122, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p115.pdf.

Peters:1979:TTS

[PW79] N. Peters and B. Watermann. Three types of skin papillomas of flatfishes and their causes. Marine Ecology Progress Series, 1:269– 276, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p269.pdf.

Rosenthal:1979:HGH

[RI79] H. Rosenthal and T. Iwai. Hatching glands in herring embryos. Marine Ecology Progress Series, 1:123–127, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/1/m001p123.pdf.

Rieper:1979:MAS

[Rie79] M. Rieper. Macro-autoradiographic studies on North Sea sediment bacteria. Marine Ecology Progress Series, 1:337–345, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p337.pdf.

Rinkevich:1979:RRSa

[RL79a] B. Rinkevich and Y. Loya. The reproduction of the Red Sea coral Stylophora pistillata. I. Gonads and planulae. Marine Ecology Progress Series, 1:133–144, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p133.pdf.

Rinkevich:1979:RRSb

[RL79b] B. Rinkevich and Y. Loya. The reproduction of the Red Sea coral Stylophora pistillata. II. Synchronization in breeding and seasonality of planulae shedding. Marine Ecology Progress Series, 1:145– 152, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p145.pdf.

Reddy:1979:EAA

[RM79] N. A. Reddy and N. R. Menon. Effects of ammonia and ammonium on tolerance and byssogenesis in Perna viridis. Marine Ecology Progress Series, 1:315–321, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/1/m001p315.pdf.

Randlov:1979:EPR

[RR79] A. Randløv and H. U. Riisgård. Efficiency of particle retention and filtration rate in four species of ascidians. Marine Ecology Progress Series, 1:55–59, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/1/m001p055.pdf.

Scheuwimmer:1979:STS

[Sch79a] A. Scheuwimmer. Sperm transfer in the sessile gastropod Serpulorbis (Prosobranchia: Vermetidae). Marine Ecology Progress Series, 1:65–70, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p065.pdf.

Schnack:1979:FCH

[Sch79b] S. B. Schnack. Feeding of Calanus helgolandicus on phytoplankton mixtures. Marine Ecology Progress Series, 1:41–47, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p041.pdf.

Schottler:1979:AMT

[Sch79c] U. Schottler. On the anaerobic metabolism of three species of Nereis (Annelida). Marine Ecology Progress Series, 1:249–254, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p249.pdf.

Spindler:1979:LPR

[SHB^{+}79] M. Spindler, C. Hernleben, U. Bayer, A. W. H. Bé, and O. R. Anderson. Lunar periodicity of reproduction in the planktonic foraminifer Hastigerina pelagica. Marine Ecology Progress Series, 1:61–64, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p061.pdf.

Sheppard:1979:IAB

[She79] C. R. C. Sheppard. Interspecific aggression between reef corals with reference to their distribution. Marine Ecology Progress Series, 1: 237–247, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p237.pdf.

Siebers:1979:TUD

[Sie79] D. Siebers. Transintegumentary uptake of dissolved amino acids in the sea star Asterias rubens. A reassessment of its nutritional role with special reference to the significance of heterotrophic bacteria. Marine Ecology Progress Series, 1:169–177, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p169.pdf.

Simmons:1979:ASD

[Sim79] G. M. Simmons, Jr. Abundance and size distribution of particulate matter fractions near a Caribbean bank barrier reef. Marine Ecology Progress Series, 1:7–11, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p007.pdf.

Spaargaren:1979:HPB

[Spa79] D. H. Spaargaren. Hydrodynamic properties of benthic marine crustacea. I. Specific gravity and drag coefficients. Marine Ecology Progress Series, 1:351–359, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/1/m001p351.pdf.

Sulkin:1979:LRB

[SPv79] S. D. Sulkin, I. Phillips, and W. van Heukelem. On the locomotory rhythm of brachyuran crab larvae and its significance in vertical migration. Marine Ecology Progress Series, 1:331–335, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p331.pdf.

Tiemann:1979:ETC

[TB79] H. Tiemann and K.-H. Betz. Elutriation: Theoretical considerations and methodological improvements. Marine Ecology Progress Series, 1:277–281, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/1/m001p277.pdf.

Thiel:1979:FQD

[Thi79] H. Thiel. First quantitative data on Red Sea deep benthos. Marine Ecology Progress Series, 1:347–350, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p347.pdf.

Theede:1979:TSE

[TSF79] H. Theede, N. Scholz, and H. Fischer. Temperature and salinity effects on the acute toxicity of cadmium to Laomedea loveni (Hydrozoa). Marine Ecology Progress Series, 1:13–19, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/1/m001p013.pdf.

Vahl:1979:ASD

[VD79] O. Vahl and J. Davenport. Apparent specific dynamic action of food in the fish Blennius pholis. Marine Ecology Progress Series, 1: 109–113, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p109.pdf.

Walker:1979:MDD

[Wal79] I. Walker. Mechanisms of density-dependent population regulation in the marine copepod Amphiascoides sp. (Harpacticoida). Marine Ecology Progress Series, 1:209–221, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/1/m001p209.pdf.

Weinberg:1979:EDS

[Wei79] J. R. Weinberg. Ecological determinants of spionid distributions within dense patches of deposit-feeding polychaete Axiothella rubrocincta. Marine Ecology Progress Series, 1:301–314, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p301.pdf.

Wolfson:1979:MLO

[WVDL79] A. Wolfson, G. Van Blaricom, N. Davis, and G. S. Lewbel. The marine life of an offshore oil platform. Marine Ecology Progress Series, 1:81–89, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/1/m001p081.pdf.

Zaika:1979:SPF

[ZM79] V. E. Zaika and N. P. Makarova. Specific production of free-living marine nematodes. Marine Ecology Progress Series, 1:153–158, 1979. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/1/
m001p153.pdf.